OpenStudy (anonymous):

integral of dx/x^3 + 4x from 1 to 2

OpenStudy (anonymous):

looks like partial fractions to me.

OpenStudy (anonymous):

integration by parts?

OpenStudy (anonymous):

i would use partial fractions. in other words write \[\frac{1}{x(x^2+4)}=\frac{1}{4x}-\frac{x}{4(x^2+4)}\]

OpenStudy (anonymous):

then the integration is easy the hard part is figuring out what i wrote. do you know how to do that?

OpenStudy (anonymous):

umm no, i do not

OpenStudy (anonymous):

have you seen partial fraction decomposition before?

OpenStudy (anonymous):

i don't think i've learned that yet, so could you do it by integration by parts?

OpenStudy (anonymous):

start by factoring the denominator. so you have \[\frac{1}{x(x^2+4)}\] then rewrite as \[\frac{a}{x}+\frac{bx+c}{x^2+4}\] and solve for \[a,b,c\]

OpenStudy (anonymous):

to be honest i am not sure how you would do this by parts, but there is more than one way to skin a cat. maybe amistre or jamesj can help with that part. it is pretty easy once you have broken it into two pieces

OpenStudy (anonymous):

okay thank you for trying!

OpenStudy (anonymous):

is this from a text, because it really looks like a set up for partial fractions

OpenStudy (amistre64):

\[\frac{d}{dx}\ tan^{-1}(\frac{x}{a})=\frac{1}{a}*\frac{1}{\frac{x^2}{a^2}+1}=\frac{1}{\frac{x^2}{a}+a}=\frac{a}{x^2+a^2}\] maybe?

OpenStudy (anonymous):

i am sticking with partial fractions. really i see no other way, and i have a feeling that although there might be one it will be extremely messy. once you have the partial fractions the integral pops right out. it is \[\frac{1}{4}\ln(x)-\frac{1}{8}\ln(x^2+4)\]

OpenStudy (amistre64):

partials are nice :)

OpenStudy (anonymous):

that is what my dentist tried to tell me

OpenStudy (amistre64):

\[\int\frac{1}{2x}*\frac{2}{(x^2+2^2)}dx\] might work for IBP

OpenStudy (amistre64):

or im over thinking it :)

OpenStudy (amistre64):

\[\int\frac{1}{x(x^2+2^2)}dx=\frac{ln(x)}{(x^2+2^2)}-\int ln(x)..dx\]lol